OpenStudy (anonymous):

If the bases are 1/2 and 1/4, and the height is 1, what is the area

OpenStudy (anonymous):

I'm guessing you're trying to find the area of a triangle?

OpenStudy (jdoe0001):

got picture?

OpenStudy (anonymous):

wouldn't you times 1/2 by 1/4 by 1 and it will equal 1/8 \[\frac{ 1 }{ 2 }\times \frac{ 1 }{ 4 }\times1=\frac{ 1 }{ 8 }\]

OpenStudy (anonymous):

no i got the awnser but thanks you guys ;)